OpenStudy (anonymous):

Two trains start at the same time from Delhi and Pune and proceed towards each other at the rate of 80 km/h"1 and 90 km/h"1 respectively. When they meet, it is found that one train has travelled 200 km more than the other. Find the distance between Delhi and Pune.

OpenStudy (raden):

(200+x)/90 = x/80 solve for x, can u ?

OpenStudy (raden):

cross product then simplify ...

OpenStudy (anonymous):

x=1600 +200=1800

OpenStudy (raden):

(200+x)/90 = x/80 90x = 80(200+x) 90x = 16,000 + 80x subtract by 80x from both sides, will giving us 90x - 80x = 16,000 + 80x - 80x 10x = 16,000 divide by 10 to both sides, 10x/10 = 16,000/10 x = 1,600 thus, the distance between Delhi and Pune = 1,600 km

OpenStudy (anonymous):

i think @luis rivera is correct

OpenStudy (anonymous):

i have pc now

OpenStudy (anonymous):

so i am logging out.sorry

OpenStudy (raden):